complete the square, and find the roots of the quadratic equation
Tutor plz help ?

x2 + 16x = 0

Answer by stanbon(75887)

x^2 + 16x = 0
x^2 + 16x + (16/2)^2 = 0 + (16/2)^2
x^2 + 16x + 64 = 64
Factor:
(x+8)^2 = 64
Take the square root of both sides:
x+8 = +8 or x+8 = -8
x = 0 or x = -16
======================
Cheers,
Stan H.